OpenStudy (anonymous):

Calculate the percentage of water sodium carbonate decahydrate Na2CO3 * 10H20

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

such percentage is: \[p = \frac{{{\text{molecular weight of }}\left( {{\text{10}}{{\text{H}}_{\text{2}}}{\text{O}}} \right)}}{{{\text{molecular weight of }}\left( {{\text{N}}{{\text{a}}_{\text{2}}}{\text{C}}{{\text{O}}_{\text{3}}} \cdot {\text{10}}{{\text{H}}_{\text{2}}}{\text{O}}} \right)}} \times 100\]